OpenStudy (juana02):

Use the distributive property to multiply. Then, if possible, simplify the resulting expression. -1(a+b)

OpenStudy (solomonzelman):

\(\large\color{black}{-1(\color{red}{a}+\color{blue}{b})~~~~~~~~~~~~~~~~~~\Rightarrow~~~~~~~~~~~~~~(-1)\times \color{red}{a} ~~~+~~~(-1)\times \color{blue}{b}}\)

OpenStudy (solomonzelman):

so how would you distribute this expression?

OpenStudy (juana02):

a+b?

OpenStudy (solomonzelman):

no, please try again.... if you don;t understand what I told you in my first reply, then say so please

OpenStudy (juana02):

i dont understand

OpenStudy (solomonzelman):

lets start from very little: when I say: \(\large\color{blue}{ab}\) do you know that I mean \(\large\color{blue}{a \times b}\) by that?

OpenStudy (solomonzelman):

just asking you, if you know that \(\large\color{blue}{ab}\) is same as \(\large\color{blue}{a \times b}\) ?

OpenStudy (juana02):

yes its a different form for same thing

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So if we have: \(\large\color{blue}{x(v+d)}\) then we distribute it to, \(\large\color{blue}{xv+xd}\)

OpenStudy (solomonzelman):

now, if you had, \(\large\color{blue}{-1(a+b)}\) , what would you then be getting?

OpenStudy (juana02):

-1a+-1b?

OpenStudy (solomonzelman):

yes, but that can be written simpler

OpenStudy (solomonzelman):

\(\large\color{blue}{-1a}\) is same as \(\large\color{blue}{-a}\). \(\large\color{blue}{+-1b}\) is same as \(\large\color{blue}{-b}\).

OpenStudy (juana02):

it will be -a+-b?

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ -a+-b }\), which is same as \(\large\color{black}{ -a-b }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ because for any 2 numbers \(\large\color{black}{ d }\) and \(\large\color{black}{ v }\) \(\large\color{black}{ d+-v=d-v }\)

OpenStudy (juana02):

ok, how do I use distributive property to simplify this one ? 24(x/3-1/8)?

OpenStudy (solomonzelman):

so you are done with the previous problem already...

OpenStudy (solomonzelman):

\(\large\color{black}{ 24 \left(\begin{matrix} \frac{\LARGE x}{\LARGE 3} -\frac{\LARGE 1}{\LARGE 8} \\ \end{matrix}\right) }\)

OpenStudy (solomonzelman):

1st find the common denominator and subtract the fractions. then multiply the fraction you got (inside the parenthesis) times 24.

OpenStudy (juana02):

yes this is the correct one

OpenStudy (solomonzelman):

okay, and what do you get when you subtract the fractions inside the parenthesis?

OpenStudy (solomonzelman):

(the least common denominator between 3 and 8, is 24)

OpenStudy (solomonzelman):

Do you need help subtraction this: \(\large\color{black}{ \frac{\LARGE x}{\LARGE 3} -\frac{\LARGE 1}{\LARGE 8} }\) ?

OpenStudy (juana02):

x/5

OpenStudy (solomonzelman):

Do you need help subtraction this: \(\large\color{black}{ \frac{\LARGE x}{\LARGE 3} -\frac{\LARGE 1}{\LARGE 8} }\) ?

OpenStudy (juana02):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE x}{\LARGE 3}- \frac{\LARGE 1}{\LARGE 8} }\) \(\large\color{black}{ \frac{\LARGE 8x}{\LARGE 24}- \frac{\LARGE 3}{\LARGE 24} }\)

OpenStudy (juana02):

oh ok u did cross mulitply

OpenStudy (solomonzelman):

multiplied: ~ top & bottom of first fraction times 8 ~ top & bottom of the second fraction times 3 (no cross multiplication)

OpenStudy (juana02):

and then u used the common denominator?

OpenStudy (solomonzelman):

yup, so what do I get?

OpenStudy (juana02):

5x?

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE x}{\LARGE 3}-\frac{\LARGE 1}{\LARGE 8}=\frac{\LARGE 8x}{\LARGE 24}-\frac{\LARGE 3}{\LARGE 24}=\frac{\LARGE 8x-3}{\LARGE 24} }\)

OpenStudy (solomonzelman):

then multiply times 24

OpenStudy (solomonzelman):

\(\large\color{black}{ 24 \times \frac{\LARGE 8x-3}{\LARGE 24} }\)

OpenStudy (solomonzelman):

\(\large\color{red}{ 24 }\)s cancel, you get?

OpenStudy (juana02):

5

OpenStudy (solomonzelman):

no

OpenStudy (juana02):

1/5